Suppose (1,1) is a critical point of a function f with continuous second derivatives. In each case, what can you say about f? (6 Pts) (a) fxx(1,1) = −3, fxy(1,1) = 1, fyy(1,1) = 2 (b) fxx(1,1) = −4, fxy(1,1) = 2, fyy(1,1) = −1 (c) fxx(1,1) = −4, fxy(1,1) = −2, fyy(1,1) = −2 (d) fxx(1,1) = 4, fxy(1,1) = 3, fyy(1,1) = 3
Answer #1

Alote: - let f(x,y) have continuous second order partial derivatives Critical point (a, b) & let D = fx. (a, b) fry (a,b) - ((b) fax (1,1)=-4, bxy (1,1)=2, fyy (1, 1) = -3 So, D =(-4) (-1) - (2)² = 4-4 = 0 5 D = 0 So, no conclusion these values. Coin